# Error Bound Analysis for the Regularized Loss of Deep Linear Neural Networks

### v1

@@ 3.1 Preliminary Results
$\mathrm{dist}^{2}(\bm{W},\mathcal{W})=\sum_{l=1}^{L}\|\bm{W}_{l}-\bm{W}_{l}^{*}\|_{F}^{2}\geq\frac{\sum_{l=1}^{L}\|\sqrt{\lambda_{l}}\bm{W}_{l}-\sqrt{\lambda_{l}}\bm{W}_{l}^{*}\|_{F}^{2}}{\lambda_{\max}}\geq\frac{1}{\lambda_{\max}}\mathrm{dist}^{2}(\bm{Z},\mathcal{W}_{G}).$ (18)
This, together with $\mathrm{dist}(\bm{W},\mathcal{W})\leq\epsilon/\sqrt{\lambda_{\max}}$, implies $\mathrm{dist}(\bm{Z},\mathcal{W}_{G})\leq\epsilon$. Using this and (16), we obtain
Next, we have
$\mathrm{dist}^{2}(\bm{Z},\mathcal{W}_{G})=\sum_{l=1}^{L}\lambda_{l}\|\bm{W}_{l}-\hat{\bm{W}}^{*}_{l}\|_{F}^{2}\geq\lambda_{\min}\sum_{l=1}^{L}\|\bm{W}_{l}-\hat{\bm{W}}^{*}_{l}\|_{F}^{2}\geq\lambda_{\min}\mathrm{dist}^{2}(\bm{W},\mathcal{W}).$ (20)
Moreover, we have for each $l\in[L]$,
@@
$=\frac{\lambda}{\sqrt{\lambda_{l}}}\nabla_{\bm{W}_{l}}F(\bm{W}).$ (21)
This, together with (19) and (20), yields
$\mathrm{dist}(\bm{W},\mathcal{W})$ $\leq\frac{1}{\sqrt{\lambda_{\min}}}\mathrm{dist}(\bm{Z},\mathcal{W}_{G})\leq\frac{\kappa}{\sqrt{\lambda_{\min}}}\|\nabla G(\bm{W})\|_{F}\leq\frac{\kappa\lambda}{\lambda_{\min}}\|\nabla F(\bm{W})\|_{F}.$
(iii) Obviously, each critical point of Problem (13) satisfies $\nabla_{\bm{W}_{l}}G(\bm{W})=0$. This, together with (5) and (14), directly implies the desired result. ∎
Using (i) and (ii) of the above lemma, it suffices to characterize the critical points and establish the error bound for Problem (13) in the rest of this section. Moreover, using (iii) of the above lemma, we assume without loss of generality that
$\bm{Y}=\mathrm{BlkDiag}\left(\tilde{\bm{\Sigma}}_{Y},\bm{0}_{(d_{L}-d_{\min})\times(d_{0}-d_{\min})}\right)\in\mathbb{R}^{d_{L}\times d_{0}},$ (22)
where $\tilde{\bm{\Sigma}}_{Y}=\mathrm{diag}(y_{1},\dots,y_{d_{\min}})$ with $y_{1}\geq y_{2}\geq\dots\geq y_{d_{\min}}\geq 0$ being singular values. According to (6), we write
$\tilde{\bm{\Sigma}}_{Y}=\mathrm{BlkDiag}\left(y_{s_{1}}\bm{I}_{h_{1}},\dots,y_{s_{p_{Y}}}\bm{I}_{h_{p_{Y}}},\bm{0}_{d_{\min}-r_{Y}}\right)\in\mathbb{R}^{d_{\min}\times d_{\min}}.$ (23)
Moreover, we define
$\delta_{y}:=\min\left\{|y_{s_{i}}-y_{s_{i+1}}|:i\in[p_{Y}]\right\}.$ (24)
Throughout this section, we will consistently use the above notation in all proofs.
3.2 Analysis of the Set of Critical Points
In this subsection, we focus on characterizing all critical points of Problem (2) explicitly by studying the critical points of Problem (13). To begin, we present a lemma to show that the weight matrices $\{\bm{W}_{l}\}_{l=1}^{L}$ at any critical point of Problem (13) are balanced.
@@ 3.2 Analysis of the Set of Critical Points
Let $(\bm{W}_{1},\dots,\bm{W}_{L})$ be a critical point of Problem (13). The following statements hold:
(i) It holds that
$\bm{W}_{l}\bm{W}_{l}^{T}=\bm{W}_{l+1}^{T}\bm{W}_{l+1},\ \forall l\in[L-1].$ (25)
(ii) It holds that
$(\bm{W}_{l}\bm{W}^{T}_{l})^{L-1}\bm{W}_{l}-\sqrt{\lambda}\bm{W}_{L:l+1}^{T}\bm{Y}\bm{W}_{l-1:1}^{T}+\lambda\bm{W}_{l}=0,\ \forall l\in[L].$ (26)
The proof of this lemma is deferred to Section 4.1. Recall the notions in (6), (22), and (23). Leveraging Lemma 1 and Lemma 2, we are ready to characterize the set of critical points (15) as follows.
Proposition 1.
Suppose that 1 holds and $\bm{Y}\in\mathbb{R}^{d_{L}\times d_{0}}$ takes the form of (22). The critical point set (15) of Problem (13) can be expressed as
$\mathcal{W}_{G}=\left\{\bm{W}:\begin{array}[]{l}\bm{\Sigma}_{l}=\mathrm{BlkDiag}\left(\mathrm{diag}(\bm{\sigma}),\bm{0}\right)\in\mathbb{R}^{d_{l}\times d_{l-1}},\ \forall l\in[L],\ (\bm{\sigma},\bm{\Pi})\in\mathcal{B},\\ \bm{W}_{1}=\bm{Q}_{2}\bm{\Sigma}_{1}\mathrm{BlkDiag}\left(\bm{\Pi},\bm{I}\right)\mathrm{BlkDiag}\left(\bm{O}_{1},\dots,\bm{O}_{p_{Y}},\bm{O}_{p_{Y}+1}\right),\\ \bm{W}_{l}=\bm{Q}_{l+1}\bm{\Sigma}_{l}\bm{Q}_{l}^{T},\ l=2,\dots,L-1,\ \bm{Q}_{l}\in\mathcal{O}^{d_{l-1}},\ l=2,\dots,L,\\ \bm{W}_{L}=\mathrm{BlkDiag}\left(\bm{O}_{1}^{T},\dots,\bm{O}_{p_{Y}}^{T},\widehat{\bm{O}}_{p_{Y}+1}^{T}\right)\mathrm{BlkDiag}\left(\bm{\Pi}^{T},\bm{I}\right)\bm{\Sigma}_{L}\bm{Q}_{L}^{T},\\ \bm{O}_{i}\in\mathcal{O}^{h_{i}},\ \forall i\in[p_{Y}],\ \bm{O}_{p_{Y}+1}\in\mathcal{O}^{d_{0}-r_{Y}},\ \widehat{\bm{O}}_{p_{Y}+1}\in\mathcal{O}^{d_{L}-r_{Y}}\end{array}\right\},$ (32)
where
$\mathcal{A}:=\left\{\bm{\sigma}\in\mathbb{R}^{d_{\min}}:\sigma_{i}^{2L-1}-\sqrt{\lambda}y_{i}\sigma_{i}^{L-1}+\lambda\sigma_{i}=0,\ \sigma_{i}\geq 0,\ \forall i\in[d_{\min}]\right\},$ (33)
:=$\mathcal{B}:=\left\{(\bm{\sigma},\bm{\Pi})\in\mathbb{R}^{d_{\min}}\times\mathcal{P}^{d_{\min}}:\bm{\sigma}=\bm{\Pi}\bm{a},\ \bm{a}\in\mathcal{A},\ \sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{d_{\min}}\right\}.$ (34)
The proof of this proposition is deferred to Section 4.1. Notably, each $\bm{W}_{l}$ in (32) is represented in an SVD form with singular values selected from $\cal A$ up to a permutation determined by $\cal B$. Based on this proposition, we can further simplify the above structure of the critical point set by removing the permutation in (32).
Theorem 3.
Suppose that 1 holds and $\bm{Y}\in\mathbb{R}^{d_{L}\times d_{0}}$ takes the form of (22). Then, the critical point set (15) of Problem (13) can be expressed as
$\mathcal{W}_{G}=\left\{\bm{W}:\begin{array}[]{l}\bm{\Sigma}_{l}=\mathrm{BlkDiag}\left(\mathrm{diag}(\bm{\sigma}),\bm{0}\right)\in\mathbb{R}^{d_{l}\times d_{l-1}},\ \forall l\in[L],\ \bm{\sigma}\in\mathcal{A},\\ \bm{W}_{1}=\bm{Q}_{2}\bm{\Sigma}_{1}\mathrm{BlkDiag}\left(\bm{O}_{1},\dots,\bm{O}_{p_{Y}},\bm{O}_{p_{Y}+1}\right),\\ \bm{W}_{l}=\bm{Q}_{l+1}\bm{\Sigma}_{l}\bm{Q}_{l}^{T},\ l=2,\dots,L-1,\ \bm{Q}_{l}\in\mathcal{O}^{d_{l-1}},\ l=2,\dots,L,\\ \bm{W}_{L}=\mathrm{BlkDiag}\left(\bm{O}_{1}^{T},\dots,\bm{O}_{p_{Y}}^{T},\widehat{\bm{O}}_{p_{Y}+1}^{T}\right)\bm{\Sigma}_{L}\bm{Q}_{L}^{T},\\ \bm{O}_{i}\in\mathcal{O}^{h_{i}},\ \forall i\in[p_{Y}],\ \bm{O}_{p_{Y}+1}\in\mathcal{O}^{d_{0}-r_{Y}},\ \widehat{\bm{O}}_{p_{Y}+1}\in\mathcal{O}^{d_{L}-r_{Y}}\end{array}\right\},$ (40)
where $\cal A$ is defined in (33).

### v3

@@ 3.2.1 Preliminary Results
$\geq\frac{\sqrt{\lambda}}{4}\min\left\{\min_{i\in[s_{p_{Y}}]}\left|\lambda-y_{i}^{2}\right|,\lambda\right\}\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{0}})$
where the third inequality follows from (7) and $\bm{Y}=\bm{\Sigma}_{Y}$ defined in (3), $\|\bm{A}\bm{B}\|_{F}\geq\sigma_{\min}(\bm{A})\|\bm{B}\|_{F}$, and $\|\bm{W}_{1}\|_{F}\geq\|\bm{W}_{2}\|_{F}$, and the fourth inequality uses $\|\bm{W}\|_{F}\leq\epsilon_{\bm{0}}$ and (31). This directly implies (30).
Next, we focus on $L\geq 3$. According to 1 and (29), we have $\mathcal{W}_{\bm{0}}=\{(\bm{0},\bm{0},\dots,\bm{0})\}$. Then, we have $\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{\bm{0}})=\sum_{l=1}^{L}\|\bm{W}_{l}\|_{F}^{2}\leq\epsilon_{\bm{0}}.$ Let $k\in[L]$ be such that $k\in\operatorname{\arg\max}_{l\in[L]}\|\bm{W}_{l}\|_{F}$. Then, we compute
@@
[…]
Thus, we have $\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{\bm{0}})\leq L\|\bm{W}_{k}\|_{F}^{2}\leq\frac{9L}{4\lambda^{2}}\left\|\nabla_{k}G(\bm{W})\right\|_{F}^{2},$ which directly implies the desired result. ∎
According to this proposition, it suffices to consider $\bm{\sigma}^{*}\neq\bm{0}$ from now on. Let $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$ be arbitrary and define
$\sigma_{\min}^{*}:=\min\left\{\sigma_{i}^{*}\neq 0:i\in[d_{\min}]\right\},\ \sigma_{\max}^{*}:=\max\left\{\sigma_{i}^{*}:i\in[d_{\min}]\right\},\ \text{and}\ \ r_{\sigma}:=\|\bm{\sigma}^{*}\|_{0}.$ (34)
By the definition of $\delta_{\sigma}$ in (28) and $\bm{0}\in\mathcal{Y}$, we have $\delta_{\sigma}\leq\sigma^{*}_{\min}$ for each $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}$. Then, we show the following results.
@@
Suppose that Assumption 1 holds. Let $\bm{W}=(\bm{W}_{1},\dots,\bm{W}_{L})$ and $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$ be arbitrary such that $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})<{\sigma^{*}_{\min}}/{2}$. Then the following statements hold:
$\text{(i)}\quad\frac{\sigma_{\max}^{*}}{2}\leq\|\bm{W}_{l}\|\leq\frac{3\sigma_{\max}^{*}}{2},\ \sigma_{i}(\bm{W}_{l})\geq\frac{\sigma_{\min}^{*}}{2},\ \forall i\in[r_{\sigma}],\ l\in[L].$ (35)
$\text{(ii)}\quad\|\bm{W}_{l+1}^{T}\bm{W}_{l+1}-\bm{W}_{l}\bm{W}_{l}^{T}\|_{F}\leq\frac{3\sqrt{2}\sigma_{\max}^{*}}{4\lambda}\|\nabla G(\bm{W})\|_{F},\ \forall l\in[L-1].$ (36)
$\text{(iii)}\quad|\sigma_{i}(\bm{W}_{l})-\sigma_{i}(\bm{W}_{l+1})|\leq\frac{3\sqrt{2}\sigma_{\max}^{*}}{4\lambda\sigma_{\min}^{*}}\|\nabla G(\bm{W})\|_{F},\ \forall i\in[r_{\sigma}],\ l\in[L-1].$ (37)
$\text{(iv)}\quad\left\|\bm{W}_{j:i}^{T}\bm{W}_{j:i}-(\bm{W}^{T}_{i}\bm{W}_{i})^{j-i+1}\right\|_{F}\leq\frac{(j-i)(j-i+1)}{2\sqrt{2}\lambda}\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2j-2i+1}\|\nabla G(\bm{W})\|_{F},\ \forall i\leq j,$ (38)
$\qquad\ \ \left\|\bm{W}_{i:j}\bm{W}_{i:j}^{T}-(\bm{W}_{i}\bm{W}_{i}^{T})^{i-j+1}\right\|_{F}\leq\frac{(i-j)(i-j+1)}{2\sqrt{2}\lambda}\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2i-2j+1}\|\nabla G(\bm{W})\|_{F},\ \forall i>j.$ (39)